# CMSO-transducing tree-like graph decompositions

### v2

@@ 5.1 Transducing weakly-bipartitive trees
By Theorem 5.4, two neighbors $x$ and $y$ of $t$ are consecutive in $<_{t}$ if and only if, the bipartition $\big{\{}L(T^{t}_{x})\cup L(T^{t}_{y}),U\setminus\big{(}L(T^{t}_{x})\cup L(T^{t}_{y})\big{)}\big{\}}$ is a member of $\mathcal{B}$ and at least one of $x$ or $y$ is not extremal in $<_{t}$. Hence, fixing, for each $\mathsf{linear}$-labeled node $t$, the two extremals of $<_{t}$ is enough to define $<_{t}$ on $\mathbb{B}\sqcup\mathbb{T^{\prime}}$ in $\operatorname{MSO}$. To this end, we consider a rooted version of $T$, in which the non-deterministically-guessed root $r$ is not labeled by $\mathsf{linear}$, and we fix the parent $p$ of each $\mathsf{linear}$-labeled node $t$ as the minimum of $<_{t}$. We also guess a set $M$ of nodes, such that in $T$ rooted in $r$, exactly one child $m$ of each $\mathsf{linear}$-labeled node $t$ with parent $p$ belongs to $M$, and furthermore, the pair $\big{\{}L(T^{t}_{p})\cup L(T^{t}_{m}),U\setminus\big{(}L(T^{t}_{p})\cup L(T^{t}_{m})\big{)}\big{\}}$ is a member of $\mathcal{B}$. Thus, $m$ can be considered as the maximum of $<_{t}$. Equipped with the relation $\mathsf{ancestor}$ (induced by the rooting of $T$ in $r$) and the colouring $M$, we are now able to define $\phi_{\mathsf{betweeness}}$ in $\operatorname{MSO}$. Indeed, for two neighbours $x,y$ of a node $t$ labeled $\mathsf{linear}$, with parent $p$, and child $m\in M$, it holds that $x<_{t}y$ if and only if there is a set $X$ containing both $L(T_{p}^{t})$ and $L(T_{x}^{t})$, excluding $L(T_{y}^{t})$ and $L(T_{m}^{t})$, and satisfying $(X,U\setminus X)\in\mathcal{B}$. Hence, we can define in $\operatorname{MSO}$, $x<_{t}y$ as follows:
$x<_{t}y:=\exists p\exists m\Big{[}\mathsf{t}\text{-}\mathsf{edge}(t,x)\land\mathsf{t}\text{-}\mathsf{edge}(t,y)\land\mathsf{parent}(p,t)\land\mathsf{t}\text{-}\mathsf{edge}(t,m)\land M(m)\land\exists X\exists Y$
$\Big{(}\mathsf{BIPART}(X,Y)\land\big{(}\mathsf{leafset}(t,p)\cup\mathsf{leafset}(t,x)\big{)}\subseteq X\land\big{(}\mathsf{leafset}(t,y)\cup\mathsf{leafset}(t,m)\big{)}\subseteq Y\Big{)}\Big{]}.$
Note that any node with a child in $M$ must necessarily be labeled $\mathsf{linear}$ and hence:
$\phi_{\mathsf{betweeness}}(t,x,y,z)$ $:=(x<_{t}y<_{t}z)\lor(z<_{t}y<_{t}x).$
@@
Indeed, considering a bipartitive family $\mathcal{B}$ makes the use of nodes labeled $\mathsf{linear}$ in the associated tree obsolete, as the following theorem states.
Theorem 5.7 (Theorem 4 in[17]).
Let $\mathcal{B}$ be a weakly-biaprtitive family and $(T,\lambda,(<_{t})_{t\in\lambda^{-1}(\mathsf{linear})})$ be the weakly-bipartitive tree it induces. If $\mathcal{B}$ is bipartitive, then $\lambda^{-1}(\mathsf{linear})=\emptyset$ and all $<_{t}$ are empty.
@@
As a consequence, we can capture any bipartitive family $\mathcal{B}$ by the simpler structure $(T,\lambda)$, called the bipartitive tree induced by $\mathcal{B}$ (or or the bipartitive tree of $\mathcal{B}$), where $\lambda$ is now a labeling function labeling each inner nodes of $T$ by either $\mathsf{degenerate}$ or $\mathsf{prime}$. Hence, we obtain the following as a corollary from Theorem 5.5, where naturally the bipartitive tree $(T,\lambda)$ of a bipartitive family $\mathcal{B}$ is modeled by a $\{\mathsf{t}\text{-}\mathsf{edge},\mathsf{degenerate}\}$-structure $\mathbb{B}$.
Corollary 5.8.
@@
There exists a non-deterministic $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction $\tau$ such that, for each bipartitive set system $(U,\mathcal{B})$ represented as the $\{\mathsf{BIPART}\}$-structure $\mathbb{B}$, $\tau(\mathbb{B})$ is non-empty and every output in $\tau(\mathbb{B})$ is equal to some $\{\mathsf{t}\text{-}\mathsf{edge},\mathsf{degenerate}\}$-structure $\mathbb{T}$ representing the bipartitive tree $(T,\mathsf{degenerate})$ of $(U,\mathcal{B})$.
5.2 Application to split decomposition
@@
Let $G$ be a directed graph. A split in $G$ is a bipartition $\{X,Y\}$ of $V(G)$ such that there are subsets $X^{\operatorname{in}},X^{\operatorname{out}}\subseteq X$, $Y^{\operatorname{in}},Y^{\operatorname{out}}\subseteq Y$ such that for all pairs of vertices $x\in X$ and $y\in Y$ the edge $xy\in E(G)$ if and only if either $x\in X^{\operatorname{out}}$ and $y\in Y^{\operatorname{in}}$ or $x\in Y^{\operatorname{out}}$ and $y\in X^{\operatorname{in}}$. Note that if $G$ is connected then $X^{\operatorname{in}}\cup X^{\operatorname{out}}\not=\emptyset$ and $Y^{\operatorname{in}}\cup Y^{\operatorname{out}}\not=\emptyset$. We call splits $\{X,Y\}$ for which either $|X|=1$ or $|Y|=1$ trivial splits. In a family $\mathcal{B}$ of splits we use the notion of strong splits to coincide with the notion of strong bipartitions.
With a split $\{X,Y\}$ in a connected graph $G$ we associate the following decomposition of $G$ into graphs $G_{X}$ and $G_{Y}$. Let $v_{X}^{Y},v_{Y}^{X}\notin V(G)$ be two new vertices, called marker vertices. Define $G_{X}$ to be the graph obtained by adding $v_{X}^{Y}$ to $G[X]$ and edge $xv_{X}^{Y}$ for every $x\in X^{\operatorname{out}}$ and edge $v_{X}^{Y}x$ for every $x\in X^{\operatorname{in}}$. The graph $G_{Y}$ is defined equivalently. Note that this decomposition captures the graph $G$ as $G$ can be obtained as the union of $G_{X}[X]$ and $G_{Y}[Y]$ with additional edges $xy$ whenever either $xv_{X}^{Y}\in V(G_{X})$ and $v_{Y}^{X}y\in V(G_{Y})$ or $xv_{Y}^{X}\in V(G_{Y})$ and $v_{X}^{Y}y\in V(G_{X})$. We can recursively decompose $G$ along splits to obtain a decomposition into graphs $G_{1},\dots,G_{k}$. We call such a decomposition of $G$ into graphs $G_{1},\dots,G_{k}$ a split decomposition if every graph $G_{i}$ is either a singleton consisting of an original (not a marker) vertex or does not have any non-trivial split. We call the graphs $G_{i}$ components of the split decomposition. To be able to reconstruct the graph, we represent a split decomposition of $G$ into graphs $G_{1},\dots,G_{k}$ by an enriched split decomposition which is a pair $(H,F)$ consisting of a directed graph $H$ with an additional (symmetric) edge relation $F$. The graph $H$ consisting of the disjoint union $G_{1},\dots,G_{k}$. The edge relation $F$ contains an (undirected) edge between every pair of markers corresponding to one step of the decomposition. For a clear distinction, we call the edges in $H$ $\mathsf{c}$-edges and the edges in $F$ $\mathsf{t}$-edges. Note that the splits used in the decomposition are in one-to-one correspondence with the $\mathsf{t}$-edges.
We require the following theorem for our transduction.
Theorem 5.9 ([16]).
@@
If $G$ is connected then the family of splits of $G$ is weakly-bipartitive.
Hence, there is a unique split decomposition of $G$, which is called the canonical split decomposition of $G$, in which we only decompose along strong splits. Note that the tree we obtain by contracting all edges of the enriched canonical split decomposition $(H,F)$ of $G$ (only leaving the t-edges) is isomorphic to the bipartitive tree of the family $\mathcal{B}$.
@@
The following (standard) technical lemma is needed both to show that a graph can be recovered from an enriched split decomposition (by an $\operatorname{MSO}$ transduction) and to define the set of edges in the enriched canonical split decomposition $(H,F)$ when transducing $(H,F)$ from $G$.
Lemma 5.10 (See e.g. Lemma 2.10 in [1]).
Let $G$ be a connected, directed graph and $(H,F)$ an enriched split decomposition of $G$. Then the following hold:
@@
  • $uv\in E(G)$ if and only if there exists a directed path from $u$ to $v$ in $(H,F)$ on which c-edges and t-edges alternate;
  • $v_{X}^{Y}v_{X^{\prime}}^{Y^{\prime}}$ is a c-edge in $(H,F)$ for vertices $v_{X}^{Y},v_{X^{\prime}}^{Y^{\prime}}$ in the same component of $H$ if and only if there is $uu^{\prime}\in E(G)$ such that $u\in Y$ and $u^{\prime}\in Y^{\prime}$.
@@
To model enriched split decompositions as relational structures we use the relational vocabulary $\{\mathsf{t}\text{-}\mathsf{edge},\mathsf{c}\text{-}\mathsf{edge}\}$ where $\mathsf{t}\text{-}\mathsf{edge}$ and $\mathsf{c}\text{-}\mathsf{edge}$ are binary relational symbols. Am enriched split decomposition $(H,F)$ of a graph $G$ is the $\{\mathsf{t}\text{-}\mathsf{edge},\mathsf{c}\text{-}\mathsf{edge}\}$-relation $\mathbb{H}$ with universe $U_{\mathbb{H}}=V(H)$, $\mathsf{t}\text{-}\mathsf{edge}_{\mathbb{H}}$ the set of t-edges in $(H,F)$ and $\mathsf{c}\text{-}\mathsf{edge}_{\mathbb{H}}$ the set of c-edges of $(H,F)$.
We use Lemma 5.2 to transduce the canonical split decomposition of a graph.
Theorem 5.11.
@@
  • let $\mathcal{B}$ be the family of splits of $G$ and $\mathbb{B}$ the $\{\mathsf{BIPART}\}$-structure modeling the weakly-bipartitive bipartition system $(V(G),\mathcal{B})$ with $U_{\mathbb{B}}=V(G)$;
  • let $T$ be the laminar tree induced by the weakly-bipartitive family $\mathcal{B}$ and let $\mathbb{T}$ be the $\{\mathsf{t}\text{-}\mathsf{edge}\}$-structure modeling it with $U_{\mathbb{T}}=V(T)$ and $L(T)=V(G)\subset U_{\mathbb{T}}$;
@@
  • $\tau_{1}$: an $\operatorname{MSO}$-interprestation which outputs the $\{\mathsf{edge},\mathsf{BIPART}\}$-structure $\mathbb{G}\sqcup\mathbb{B}$ from $\mathbb{G}$;
  • $\tau_{2}$: the non-deterministic $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction from Lemma 5.2 which outputs the $\{\mathsf{edge},\mathsf{BIPART},\mathsf{t}\text{-}\mathsf{edge}\}$-structure $\mathbb{G}\sqcup\mathbb{B}\sqcup\mathbb{T}$ on input $\mathbb{G}\sqcup\mathbb{S}$;
@@
To define $\tau_{1}$ it is sufficient to observe that there is an $\operatorname{MSO}$-formula $\phi_{\mathsf{BIPART}}(X,Y)$ with two monadic free-variables $X$ and $Y$ such that $\phi_{\mathsf{BIPART}}$ is satisfied exactly when $\{X,Y\}$ is a split in $G$. As $\tau_{2}$ is constructed in Lemma 5.2, we are left with describing how to obtain $\tau_{3}$. We first need to find unique representative for the vertices in $V(H)$. Recall that $V(H)$ consists of two marker vertices $v_{X}^{Y}$ and $v_{Y}^{X}$ for every strong bipartition $\{X,Y\}\in\mathcal{B}$. For ease of notation, we use the name $v^{V(G)\setminus\{u\}}_{\{u\}}$ for the leaf $u\in V(G)$ of $T$. Additionally, we set $v_{s}^{t}:=v_{X}^{Y}$ and $v_{t}^{s}:=v_{Y}^{X}$ where $\{s,t\}\in E(T)$ is the edge for which $L(T_{s})=X$ and $L(T_{t})=Y$ where $T_{s}$ is the component of $T-e$ containing $s$ and $T_{t}$ the component containing $t$. We use the following atomic transductions to ensure unique representatives for all vertices in $V(H)$. We first guess a colouring $R$ which identifies one non-leaf element of $\mathbb{G}\sqcup\mathbb{B}\sqcup\mathbb{T}$ as root. We use this root element to interpret an auxiliary binary relation $\mathsf{parent}$ in the usual way. Now we copy the structure $\mathbb{G}\sqcup\mathbb{B}\sqcup\mathbb{T}$ once introducing the binary relation $\mathsf{copy}_{1}$ in which $\mathsf{copy}_{1}(x,y)$ indicates that $x$ is the copy of $y$. Assume that $\{s,t\}$ is an edge in $T$ corresponding to a strong bipartition in $\mathcal{B}$ and $\mathsf{parent}(s,t)$ is satisfied. We use the original node $t$ to represent the marker $v_{t}^{s}$ and we use the copy of $t$ to represent the marker $v_{s}^{t}$. As every node has a unique parent in a rooted tree, for each marker the chosen representatives are unique. To express this we use the predicate $\mathsf{marker}(x,y,z)$ which expresses that $x$ is the representative of the marker $v_{y}^{z}$ defined as follows:
$\mathsf{marker}(x,y,z):=\Big{(}\mathsf{parent}(z,y)\land x=y\Big{)}\lor\Big{(}\mathsf{parent}(y,z)\land\mathsf{copy}_{1}(x,z)\Big{)}.$
@@
We now filter the universe only keeping all representatives of markers (recall that markers $v_{\{v\}}^{V(G)\setminus\{v\}}$ are the original vertices of $G$). We are left with defining the two edge relations $\mathsf{t}\text{-}\mathsf{edge}$ and $\mathsf{c}\text{-}\mathsf{edge}$. To do so, we use the previously defined functional predicate $\mathsf{leafset}(t,s)$.
@@
First note that markers $v_{t}^{s}$ and $v_{t^{\prime}}^{s^{\prime}}$ are in the same component of $H$ if and only if $t=t^{\prime}$. By Lemma 5.10 $v_{t}^{s}v_{t}^{s^{\prime}}$ is an c-edge in $(H,F)$ if and only if there is an edge $uu^{\prime}\in E(G)$ with $u\in L(T^{t}_{s})$ and $u^{\prime}\in L(T^{t}_{s^{\prime}})$. Consequently, we can express $\mathsf{c}\text{-}\mathsf{edge}$ as follows:
$\phi_{\mathsf{c}\text{-}\mathsf{edge}}(x,x^{\prime}):=$ $\exists s\exists s^{\prime}\exists t\exists y\exists y^{\prime}\Big{(}\mathsf{t}\text{-}\mathsf{edge}(t,s)\land\mathsf{t}\text{-}\mathsf{edge}(t,s^{\prime})\land\mathsf{marker}(x,t,s)\land$
$\mathsf{marker}(x^{\prime},t,s^{\prime})\land y\in\mathsf{leafset}(t,s)\land y^{\prime}\in\mathsf{leafset}(t,s^{\prime})\land E(y,y^{\prime})\Big{)}.$
@@
Finally, the transduction only keeps the relation $\mathsf{t}\text{-}\mathsf{edge}$ and $\mathsf{c}\text{-}\mathsf{edge}$.
5.3 Application to bi-join decomposition

### v4

@@ 5.1. Transducing weakly-bipartitive trees
By 5.2, a set $\mathcal{C}$ of neighbours of $t$ is consecutive in $_{t}$ (understood as a cyclic order) if and only if, the bipartition $\big{\{}\bigcup_{c\in\mathcal{C}}L(T^{t}_{c}),U\setminus\big{(}\bigcup_{c\in\mathcal{C}}L(T^{t}_{c})\big{)}\big{\}}$ is a member of $\mathcal{B}$. Furthermore, we can express that $w<_{t}x<_{t}y<_{t}z$ in $_{t}$ (considered as a cyclic order) if there is a $_{t}-$interval $\mathcal{C}$ which does not contain $z$, but contains two $_{t}-$intervals $\mathcal{C}_{1},\mathcal{C}_{2}$ such that $w,x\in\mathcal{C}_{1}$, $y\notin\mathcal{C}_{1}$ and $w\notin\mathcal{C}_{2}$ and $x,y\in\mathcal{C}_{2}$. Hence, we can define in $\operatorname{MSO}$ the predicate $\mathsf{cross}$ as follows:
$\phi_{\mathsf{cross}}(t,x_{1},x_{2},x_{3},x_{4}):=\Big{[}\lnot\mathsf{degenerate}(t)\land\bigwedge_{i<j}x_{i}\not=x_{j}\land\bigwedge_{i}\mathsf{t}\text{-}\mathsf{edge}(t,x_{i})\land\exists X,Y,Z$
$\Big{(}\mathsf{BIPART}(X)\land\mathsf{BIPART}(Y)\land\mathsf{BIPART}(Z)\land Y\subseteq X\land Z\subseteq X\land\mathsf{leafset}(t,x_{1})\subseteq Y\setminus Z\land$
$\mathsf{leafset}(t,x_{2})\subseteq Y\cap Z\land\mathsf{leafset}(t,x_{3})\subseteq Z\setminus Y\land\mathsf{leafset}(t,x_{4})\subseteq U\setminus X\Big{)}\Big{]}.$
Note that any node which is not labeled $\mathsf{degenerate}$ and for which there is a set $\mathcal{C}$ of at least two neighbours of $t$ for which $\big{\{}\bigcup_{c\in\mathcal{C}}L(T^{t}_{c}),U\setminus\big{(}\bigcup_{c\in\mathcal{C}}L(T^{t}_{c})\big{)}\big{\}}$ is a member of $\mathcal{B}$ must be labeled $\mathsf{linear}$. Additionally, note that any node with less than three children which is not labeled $\mathsf{prime}$ must be labeled $\mathsf{degenerate}$.
@@
Indeed, considering a bipartitive family $\mathcal{B}$ makes the use of nodes labeled $\mathsf{linear}$ in the associated tree obsolete, as the following theorem states. {thmC}[[dM03, Theorem 4]] Let $\mathcal{B}$ be a weakly-biaprtitive family and let $(T,\lambda,(<_{t})_{t\in\lambda^{-1}(\mathsf{linear})})$ be the weakly-bipartitive tree it induces. If $\mathcal{B}$ is bipartitive, then $\lambda^{-1}(\mathsf{linear})=\emptyset$ and all $_{t}$ are empty. As a consequence, we can capture any bipartitive family $\mathcal{B}$ by the simpler structure $(T,\lambda)$, called the bipartitive tree induced by $\mathcal{B}$ (or the bipartitive tree of $\mathcal{B}$), where $\lambda$ is now a labeling function labeling each inner nodes of $T$ by either $\mathsf{degenerate}$ or $\mathsf{prime}$. Hence, we obtain the following as a corollary from 22, where naturally the bipartitive tree $(T,\lambda)$ of a bipartitive family $\mathcal{B}$ is modeled by a $\{\mathsf{t}\text{-}\mathsf{edge},\mathsf{degenerate}\}$-structure $\mathbb{B}$.
Corollary 23.
@@
There exists a non-deterministic $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction $\tau$ such that, for each bipartitive set system $(U,\mathcal{B})$ represented as the $\{\mathsf{BIPART}\}$-structure $\mathbb{B}$, $\tau(\mathbb{B})$ is non-empty and every output in $\tau(\mathbb{B})$ is equal to some $\{\mathsf{t}\text{-}\mathsf{edge},\mathsf{degenerate}\}$-structure $\mathbb{T}$ representing the bipartitive tree $(T,\lambda)$ of $(U,\mathcal{B})$.
5.2. Application to split decomposition
@@
Let $G$ be a directed graph. A split in $G$ is a bipartition $\{X,Y\}$ of $V(G)$ such that there are subsets $X^{\operatorname{in}},X^{\operatorname{out}}\subseteq X$, $Y^{\operatorname{in}},Y^{\operatorname{out}}\subseteq Y$ such that for all pairs of vertices $x\in X$ and $y\in Y$ the edge $xy\in E(G)$ if and only if either $x\in X^{\operatorname{out}}$ and $y\in Y^{\operatorname{in}}$ or $x\in Y^{\operatorname{out}}$ and $y\in X^{\operatorname{in}}$. Note that if $G$ is connected then $X^{\operatorname{in}}\cup X^{\operatorname{out}}\not$$\emptyset$ and $Y^{\operatorname{in}}\cup Y^{\operatorname{out}}\not$$\emptyset$. We call splits $\{X,Y\}$ for which either $|X|=1$ or $|Y|=1$ trivial splits. A graph is considered prime with respect to splits if it has only trivial splits. In a family $\mathcal{B}$ of splits we use the notion of strong splits to coincide with the notion of strong bipartitions. See Figure 8 (left) for an example of a strong split in a directed graph.
One can define split decomopositions in greater generality, however, we only require the canonical split decomposition obtained by considering only strong splits in the following. Hence, consider a directed graph $G$ and let $\mathcal{B}_{\mathsf{split}}^{G}$ be the set of all splits of $G$. The following theorem shows that $\mathcal{B}_{\mathsf{split}}^{G}$ is weakly-bipartitive.
{thmC}
@@
[[Cun82]] For every connected graph $G$ family of splits $\mathcal{B}_{\mathsf{split}}^{G}$ is weakly-bipartitive.
[…]
  • $H$ is a directed graph consisting of the disjoint union of the graphs $G_{u}$ for every inner node $u$ of $T$ and
  • $F$ is a symmetric edge relation containing edges $\vec{uv}\vec{vu}$ for every edge $uv$ of $T$.
For a clear distinction, we call the edges in $H$ $\mathsf{c}$-edges and the edges in $F$ $\mathsf{t}$-edges. Note that $\mathsf{t}$-edges correspond to the edges of $T$. See Figure 8 for an illustration.
Figure 8. A directed graph $G$ with 33 strong splits (left). The weakly-partitive tree of $G$ (middle) and the split decomposition of $G$ where $\mathsf{c}$-edges are the thin, black edges, $\mathsf{t}$-edges are the fat, dashed, blue edges and the (non-singleton) components of the decomposition are signified by grey circles.
@@
The following (standard) technical lemma is needed to show that a graph can be recovered from an enriched split decomposition (by an $\operatorname{MSO}$ transduction).
Lemma 24 (See e.g. Lemma 2.10 in [AKK17]).
@@
Let $G$ be a connected, directed graph and $(H,F)$ the enriched split decomposition of $G$. Then $uv\in E(G)$ if and only if there exists a directed path from $u$ to $v$ in $(H,F)$ on which c-edges and t-edges alternate.
@@
To model enriched split decompositions as relational structures we use the relational vocabulary $\{\mathsf{t}\text{-}\mathsf{edge},\mathsf{c}\text{-}\mathsf{edge}\}$ where $\mathsf{t}\text{-}\mathsf{edge}$ and $\mathsf{c}\text{-}\mathsf{edge}$ are binary relational symbols. An enriched split decomposition $(H,F)$ of a graph $G$ is the $\{\mathsf{t}\text{-}\mathsf{edge},\mathsf{c}\text{-}\mathsf{edge}\}$-relation $\mathbb{H}$ with universe $U_{\mathbb{H}}=V(H)$, $\mathsf{t}\text{-}\mathsf{edge}_{\mathbb{H}}$ the set of t-edges in $(H,F)$ and $\mathsf{c}\text{-}\mathsf{edge}_{\mathbb{H}}$ the set of c-edges of $(H,F)$.
Since the nodes of $T$ correspond to the graphs $G_{u}$ which are the connected components of the enriched split decomposition $(H,F)$ after removing $F$ and the vertices of $G$ are the leaves of $T$, it is clear that the 24 allows to $\operatorname{MSO}$-transduce the graph $G$ from its enriched split decomposition.
We use 21 to transduce the canonical split decomposition of a graph.
Theorem 25.
@@
  • let $\mathcal{B}_{\mathsf{split}}^{G}$ be the family of splits of $G$ and $\mathbb{B}$ the $\{\mathsf{BIPART}\}$-structure modeling the weakly-bipartitive bipartition system $(V(G),\mathcal{B}_{\mathsf{split}}^{G})$ with $U_{\mathbb{B}}=V(G)$;
  • let $T$ be the laminar tree induced by the weakly-bipartitive family $\mathcal{B}_{\mathsf{split}}^{G}$ and let $\mathbb{T}$ be the $\{\mathsf{t}\text{-}\mathsf{edge}\}$-structure modeling it with $U_{\mathbb{T}}=V(T)$ and $L(T)=V(G)\subset U_{\mathbb{T}}$;
@@
  • $\tau_{1}$: an $\operatorname{MSO}$-interpretation which outputs the $\{\mathsf{edge},\mathsf{BIPART}\}$-structure $\mathbb{G}\sqcup\mathbb{B}$ from $\mathbb{G}$;
  • $\tau_{2}$: the non-deterministic $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction from 21 which outputs the $\{\mathsf{edge},\mathsf{BIPART},\mathsf{t}\text{-}\mathsf{edge}\}$-structure $\mathbb{G}\sqcup\mathbb{B}\sqcup\mathbb{T}$ on input $\mathbb{G}\sqcup\mathbb{B}$;
@@
To define $\tau_{1}$ it is sufficient to observe that there is an $\operatorname{MSO}$-formula $\phi_{\mathsf{BIPART}}(X)$ with one monadic free-variable $X$ such that $\phi_{\mathsf{BIPART}}$ is satisfied exactly when $\{X,U\setminus X\}$ is a split in $G$. As $\tau_{2}$ is constructed in 21, we are left with describing how to obtain $\tau_{3}$. We first need to find unique representatives for the vertices in $V(H)$. Recall that $V(H)$ consists of two marker vertices $\vec{uv},\vec{vu}$ for every edge $uv$ in $T$. We use the following atomic transductions to ensure unique representatives for all vertices in $V(H)$. We first guess a colouring $R$ which identifies one non-leaf element of $\mathbb{G}\sqcup\mathbb{B}\sqcup\mathbb{T}$ as root. We use this root element to interpret an auxiliary binary relation $\mathsf{parent}$ in the usual way. Now we copy the structure $\mathbb{G}\sqcup\mathbb{B}\sqcup\mathbb{T}$ once introducing the binary relation $\mathsf{copy}_{1}$ in which $\mathsf{copy}_{1}(x,y)$ indicates that $x$ is the copy of $y$. Assume that $\{s,t\}$ is an edge in $T$ corresponding to a strong bipartition in $\mathcal{B}_{\mathsf{split}}^{G}$ and $\mathsf{parent}(s,t)$ is satisfied. We use the original node $t$ to represent the marker $\vec{st}$ and we use the copy of $t$ to represent the marker $\vec{ts}$. As every node has a unique parent in a rooted tree, for each marker the chosen representatives are unique. To express this we use the predicate $\mathsf{marker}(x,y,z)$ which expresses that $x$ is the representative of the marker $\vec{yz}$ defined as follows:
$\mathsf{marker}(x,y,z):=\Big{(}\mathsf{parent}(y,z)\land x=z\Big{)}\lor\Big{(}\mathsf{parent}(z,y)\land\mathsf{copy}_{1}(x,y)\Big{)}.$
@@
We now filter the universe only keeping all representatives of markers (recall that every original vertex $u$ of $G$ has a marker $\vec{uv}$, where $v$ is the neighbour of $u$ in $T$, representing it). We are left with defining the two edge relations $\mathsf{t}\text{-}\mathsf{edge}$ and $\mathsf{c}\text{-}\mathsf{edge}$. To do so, we use the previously defined functional predicate $\mathsf{leafset}(t,s)$ which is interpreted by the set $L(T_{s}^{t})$.
@@
First note that markers $\vec{st}$ and $\vec{s^{\prime}t^{\prime}}$ are in the same component of $H$ if and only if $s=s^{\prime}$. By definition $\vec{st}\vec{st^{\prime}}$ is a c-edge in $(H,F)$ if and only if there is an edge $uu^{\prime}\in E(G)$ with $u\in L(T^{s}_{t})$ and $u^{\prime}\in L(T^{s}_{t^{\prime}})$. Consequently, we can express $\mathsf{c}\text{-}\mathsf{edge}$ as follows:
$\phi_{\mathsf{c}\text{-}\mathsf{edge}}(x,x^{\prime}):=$ $\exists t\exists t^{\prime}\exists s\exists y\exists y^{\prime}\Big{(}\mathsf{t}\text{-}\mathsf{edge}(s,t)\land\mathsf{t}\text{-}\mathsf{edge}(s,t^{\prime})\land\mathsf{marker}(x,s,t)\land$
$\mathsf{marker}(x^{\prime},s,t^{\prime})\land y\in\mathsf{leafset}(s,t)\land y^{\prime}\in\mathsf{leafset}(s,t^{\prime})\land E(y,y^{\prime})\Big{)}.$
@@
Finally, the transduction only keeps the relation $\mathsf{t}\text{-}\mathsf{edge}$ and $\mathsf{c}\text{-}\mathsf{edge}$. We remark that any choice of $R$ yields a valid output and hence we do not need to apply filtering.
5.3. Application to bi-join decomposition